# DRAFT- Rotation Curve Fitting Model

### v2

@@ 2.5 Geometric simplifications
Population synthesis models commonly assume spherical symmetry for the stellar bulge and gas halo, but axial symmetry for the stellar disk (Schwarzschild, 1954; Freeman, 1970). However, it is a common calculational tool to use spherical symmetry for the entire integrated potential of the luminous mass distribution (Bekenstein, 2004; McGaugh, 2008; Mistele et al., 2022) because numerical integration of the thin disk is computationally intensive, requiring assumptions of underconstrained boundary conditions and relevant physical scales, which therefore add extra free parameters to the problem (Huré & Hersant, 2011). We use the Schwarzschild metric as it is the simplest spherically symmetric representation, which captures the relevant physics of the problem without excess computations. A more detailed analysis could use the interior Schwarzschild metric. We have previously tested the Kerr metric in this context (Cisneros et al., 2015), and found the difference from Schwarzschild predictions are below the observational limits.
Gravitational potentials calculated in a spherically symmetric geometry converge to those in an exponential disk geometry at lengths greater than one-third of the exponential scale-length , ie. (Chatterjee, 1987). Since this is the region where dark matter effects becomes important (van Albada et al., 1985), for proof of concept this calculational technique is acceptable. However, in the region from to , where the RCFM begins to compare galaxies, the spherical assumption overestimates the potential by a factor of . This artificat can be seen in the RCFM results in Tables 6 and 11. Implementation of a thin disk geometry would resolve this artifact of the analysis (de Vaucouleurs, 1959),(Freeman, 1970), though at significant computational expense.
@@ 3.2 Milky Way Luminous Mass Models
The two MW models differ markedly in the inner kpc; as the McGaugh MW has a triaxial bar-bulge and the Sofue MW has a de Vaucouleurs bulge. The Sofue MW model is used from to kpc, and extended with the Xue MW model from to kpc Xue et al. (2008). The Sofue and Xue MW models come from the same data and dark matter model constraints. The McGaugh Milky Way model covers from to kpc, and comes from data and MOND model constraints.
From the recent Gaia data release (DR3), Jiao, Yongjun et al. (2023) find a Keplerian decline of the rotation curve of the Milky Way in the range of to kpc, consistent with the RCFM heuristic picture. The ESA Gaia mission, taking measurements from the second Lagrange point, has revolutionized the science of the MW with unprecedented detail, statistical accuracy, and a drastic reduction in systematic uncertainties. We have replaced the McGaugh and Sofue-Xue Milky Way model velocities in the range from to kpc with the Jiao Milky Way velocities, and then shifted the two sides of the given MW models by a constant amount to match the endpoints of the Jiao velocities. See Table 1 for details.
@@ 4.1 Fitting procedure
To fit galaxies in the SPARC sample with the RCFM model, the fitting procedure is as follows. First, a Milky Way model is chosen (Sec. 3.2) and the data read in as for a series of measurements in radii. The galactic gravitational potential is computed as described (see Sec. 2.4), by numerically integrating
@@ 4.2 Evaluating goodness-of-fits
There are two metrics by which different rotation curve fitting models are compared, the resulting reduced values and the mass-to-light ratios. Since error estimates on RC velocities have not been standardized across the field (de Blok et al., 2008b; Gentile et al., 2011), values can only be compared when fitted to the same RC data. In Table 11 we compare values from three models; dark matter halo, Radial-Acceleration-Relation (RAR) (McGaugh et al., 2016; Lelli et al., 2017; Li et al., 2018) and RCFM, fitted to the SPARC database of galaxy RCs(Lelli et al., 2016) . RCFM values are remarkably low, providing confidence in the faithfulness of the model to galaxy RC data.
The average mass-to-light ratios from RCFM fits yield bulge mass-to-light ratios which are within the error estimates on population synthesis models (PSM). The RCFM disk mass-to-light ratios are approximately a factor of two larger than PSM, consistent with the artifact introduced from the spherical integration technique, see Sec. 2.5 for a full explanation. This artifact can be removed by including an exponential disk integration.
@@ 4.3 Comparing Milky Way models
The RCFM fits require a static choice for the Milky Way baryon distribution. In Table 2 and 10 we compare the RCFM fit results for the two different Milky Way model assumptions used in this paper. As can be seen in the table, for the 36 most reliable galaxies in the SPARC sample the two Milky Way models are equivalent (see Sec. 4.5). However, for the whole SPARC sample of 175 galaxies the RCFM fits using the Sofue-Xue-Jiao Milky Way performs better with an average , versus those from McGaugh-Jiao Milky Way . Based on this, values in Table 11 and figures are from RCFM fits assuming the Sofue-Xue-Jiao Milky Way.
The residuals of fits to the SPARC sample are also used to compare the two different MW models (see Sec. 3.2). Histograms of residuals normalized by the error in velocity observations are shown in Fig. 2. In all cases, residuals of model fits to observed velocity data followed a narrow distribution centered at zero with a range of km/s, albeit with heavy tail features. The behavior of the residuals did not vary greatly between MW models, suggesting that the fitting parameters in our model are robust with respect to differing MW model assumptions at this present level of analysis. With a finer grained analysis and a larger sample, the degeneracy between MW models may be resolved.
A Gaussian fit on residuals from fits using the McGaugh-Jiao MW model gave a mean of km/s, standard deviation of km/s; whereas a Gaussian fit to the residuals from fits using the Sofue-Xue-Jiao MW model gave a mean of km/s, standard deviation of km/s. The residuals are shown in Fig. 2. The small values associated with these quantities in both cases provide confidence that our fits match data closely.
@@ 4.4.1 NGC 2841
NGC 2841 is a star dominated, flocculent spiral galaxy, which historically has been regarded as a problematic case for MOND ((Gentile et al., 2011)). RAR finds a good fit for this galaxy after adjusting the Cepheid distance of Mpc, by , to Mpc. Cepheid variable stars are the most accurate distance indicators currently available. The RCFM fit is shown in Fig. 4 at the Cepheid based distance.
| Model | distance(Mpc) | |||
| Dark Matter Iso | 0.60 | 0.66 | 1.58 | 14.1 |
| RAR | 0.81 | 0.93 | 1.52 | 15.5 |
@@ 4.4.2 IC 2574
IC 2574 is a dwarf spiral galaxy, dominated by gas. RAR finds a good fit to this galaxy but adjusts the distance and inclination by to . This galaxy is problematic for dark matter model fits, as the model overestimates the inner RC out to 10 kpc (Navarro et al., 2017). RCFM fits this galaxy successfully, at the reported tip of the red giant branch distance of Mpc, Fig.5.
| Model | |||
| Dark Matter Iso | 0.77 | – | 2.51 |
| RAR | 0.07 | – | 1.44 |
@@ 4.4.3 NGC 3198
NGC 3198 is a barred spiral in Ursa Major, which has been well studied. It has been considered a problem galaxy for MOND, when the distance is a free parameter in fits (Gentile et al., 2011), as the preferred MOND distance is 2 different from that reported from Cepheids. However, in the SPARC database the RAR fit to this galaxy at the Cepheid distance reproduces the RC well. See Fig. 9 for the RCFM fit.
| Model | distance (MPC) | |||
| Dark Matter Iso | 0.52 | – | 1.31 | 13.7 |
| RAR | 0.77 | – | 2.06 | 13.8 |
@@ 4.4.4 NGC 7814 and NGC 891
| Model | |||
| Dark Matter Iso | 0.68 | 0.71 | 0.39 |
| RAR | 1.17 | 0.52 | 1.334 |
@@
| Model | |||
| Dark Matter Iso | 0.77 | 1.63 | 1.30 |
| RAR | 0.32 | — | 25.16 |
@@ 4.4.5 NGC 5055
| Model | |||
| Dark matter - Isothermal Halo | 0.26 | - | 6.19 |
| RAR | 0.56 | - | 7.42 |
@@
The second RC for NGC 5055 is from the THINGS (de Blok et al., 2008b) at a distance of Mpc from the Hubble Flow, including a stellar bulge component. RCFM fits to both sets of RC data can be seen in Fig.7, at reported distances, using the reported luminous mass models. This galaxy serves as an example of the under-constrained nature of galaxy mass modeling (Conroy et al., 2009).
| Model | |||
| Dark Matter Iso | 0.79 | 0.11 | 8.13 |
| Dark Matter NFW | 0.79 | 0.11 | 10.31 |
| RAR | 0.43 | 0.46 | 2.63 |
@@ 4.5 Free Parameter Correlation
We test a correlation between the model’s free parameter and a ratio of total luminosity by the half-light radius (effective radius). To test this correlation, we first create a subset of galaxies from the SPARC dataset by the following selection criteria:
@@
Select galaxies with the most accurate and reliable distance estimates (tip of the red giant branch, and Cepheid variable stars), rejecting all other galaxies.
@@
Exclude galaxies with quality factor as not suited to dynamical studies due to asymmetries, non-circular motions, and/or offsets between stars and gas, assigned SPARC factors as in (Lelli et al., 2016).
By this process, a training set of galaxies are selected, as reported in Table 2. We then plot the subset’s values versus the ratio of , and fit the distribution with a power law as shown in Fig. 6. The total luminosity and effective radii numbers used are as reported in the SPARC database. Luminosity measurements are taken in the wavelength of , assuming a solar absolute magnitude of 3.24 at (Oh et al., 2008) in units of solar .
| Model | |||
| RAR | 0.46 | 0.74 | 3.89 |
| RCFM (MW: Sofue-Xue-Jiao) | 1.04 | 0.87 | 1.60 |
| RCFM (MW: McGaugh-Jiao) | 1.00 | 0.87 | 1.65 |
@@
4.6 Galaxies for which model fits fail
Quality flag are assigned in the SPARC galaxy database, using the following scheme: for galaxies with high-quality HI data or hybrid H/HI rotation curves; for galaxies with minor asymmetries and/or HI data of lower quality; Q = 3 for galaxies with major asymmetries, strong non-circular motions, and/or offsets between HI and stellar distributions. They state galaxies with are not suited for detailed dynamical studies. Of the three models compared in this paper for the 175 SPARC galaxies;
The RAR fits all galaxies;
The RCFM fits 173 galaxies; fits fail for UGC06628 (), and F561-1, UGC04305 (),
The dark matter model fits 166 galaxies; fits fail for D512-2, UGC00634, NGC6789, UGC00891, UGC02023, UGC05999, UGC07232 and UGC09992 (), and F567-2 and F574-2 ().
5 Conclusions
At this time, there are many discrepancies between cosmology theory and observations which are resolved by dark matter models (Sanders, 2010; Tully et al., 2014; Naidu et al., 2022). In this paper we address only one, the flat-rotation curve problem of spiral galaxies. This choice is due to the clear symmetry presented in the rotation curve data. Other dark matter problems require cosmological model assumptions, beyond the scope of the current paper.
The rotation curve fitting model (RCFM) presented here reproduces the fitting successes of MOND and dark matter models on a sample of 175 well studied galaxy rotation curves(Lelli et al., 2016), but does not modify classical physics. We emphasize that the only input to the RCFM is the luminous mass. Independent of the interpretation of the various quantities appearing in the formulae presented in this paper, this is a one parameter fit to the data. What is more, the RCFM free parameter is highly correlated with a ratio of observable photometric parameters. Our results are presented in this paper in comparison to dark matter and RAR (Radial Acceleration Relation from a MONDian basis) results for the same galaxies in the SPARC sample of 175 rotation curves (Lelli et al., 2016). Consistent with this model paradigm, the recent Gaia DR3 data release Jiao, Yongjun et al. (2023) has shown that the rotation curve of the Milky Way demonstrates a Keplerian decline from - kpc. The upcoming Large Survey of Space and Time at the Vera C. Rubin Observatory (Željko Ivezić et al., 2019) can falsify this model if the Milky Way rotation curve beyond kpc is found to require a dark matter halo.
| Galaxy | MW: M.J.333Milky Way:M.J. is McGaugh-Jiao | MW: S.X.J.444 Milky Way:S.X.J. is Sofue-Xue-Jiao | Distance | Distance | Inc |
| Name | (Mpc) | Method 555 Distance method: 2 = tip of the red giant branch, 3 = Cepheids. Distance and inclination information from (Lelli et al., 2016). | (deg) | ||

### v4

@@ 2.5 Geometric simplifications
Population synthesis models commonly assume spherical symmetry for the stellar bulge and gas halo, but axial symmetry for the stellar disk (Schwarzschild, 1954; Freeman, 1970). However, it is a common calculational tool to use spherical symmetry for the entire integrated potential of the luminous mass distribution (Bekenstein, 2004; McGaugh, 2008; Mistele et al., 2022) because numerical integration of the thin disk is computationally intensive, requiring assumptions of under-constrained boundary conditions and relevant physical scales, which therefore add extra free parameters to the problem (Huré & Hersant, 2011). We use the spherically symmetric Schwarzschild metric here as proof of concept, as it is the simplest representation which captures the relevant physics of the problem without excess computations.
For a given mass, the gravitational potentials calculated in a spherically symmetric geometry converge to those for an exponential disk geometry at lengths greater than one-third of the exponential scale-length , ie. (Chatterjee, 1987). Since this is the region where dark matter effects becomes important (van Albada et al., 1985), this calculational technique captures the relevant physics. However, in the region where the RCFM integrations begin galaxy comparisons ( to ) the spherical assumption overestimates the potential by a factor of , which can be seen in RCFM disk mass-to-light ratio results. Implementation of a thin disk geometry would resolve this problem (de Vaucouleurs, 1959),(Freeman, 1970), though at significant computational expense.
@@ 3.2 Milky Way Luminous Mass Models
The two MW models differ markedly in the inner kpc; as the McGaugh MW has a triaxial bar-bulge and the Sofue MW has a de Vaucouleurs bulge. The Sofue MW model is used from to kpc, and extended with the Xue MW model from to kpc (Xue et al., 2008). The Sofue and Xue MW models come from the same data and a dark matter model is used for the extended rotation curve. The McGaugh Milky Way model covers from to kpc, comes from data, and the MOND model is used for the extended rotation curve.
The recent ESA Gaia data release (DR3), Jiao, Yongjun et al. (2023) demonstrates a Keplerian decline of the rotation curve of the Milky Way in the range of to kpc, consistent with this RCFM heuristic paradigm. The ESA Gaia mission, taking measurements from the second Lagrange point, has revolutionized the science of the MW with unprecedented detail, statistical accuracy, and a drastic reduction in systematic uncertainties. Since this is the most recent and best data for the MW, we replace the model Milky Way model velocities in the range from to kpc with those from Jiao, Yongjun et al. (2023). We then shift the two remaining sides of the given MW model by a constant amount to match the endpoints of the Jiao rotation curve. See Table 1 for summary details.
| Author | scale length | Model for the | Range |
| bulge/bar | extended RC | ||
| Sofue-Xue | 0.5 kpc | NFW (Navarro et al., 1996) | [0,60] kpc |
| Sofue (2013); Xue et al. (2008) | bulge | dark matter | |
| McGaugh | 2.0 kpc | MOND | [0,150] kpc |
| McGaugh (2008) | bar | (Milgrom, 1983) |
@@ 4.1 Fitting procedure
To fit galaxies in the SPARC sample with the RCFM model, the fitting procedure is as follows. First, a Milky Way baryon model is selected (Sec. 3.2) and the model data in is read in for a series of measurements in radii. The galactic gravitational potential is calculated by numerically integratingas (see Sec. 2.4),
@@ 4.2 Evaluating goodness-of-fits
There are two metrics by which different rotation curve fitting models are compared, the resulting reduced values and the mass-to-light ratios. Since error estimates on RC velocities have not been standardized across the field (de Blok et al., 2008a; Gentile et al., 2011), values can only be compared when fitted to the same RC data. In Table LABEL:table:M2Light we compare values from three models: dark matter halo, Radial-Acceleration-Relation (RAR) (McGaugh et al., 2016; Lelli et al., 2017; Li et al., 2018) and RCFM, fitted to the SPARC database of galaxy RCs (Lelli et al., 2016) . RCFM values are remarkably low, providing confidence in the faithfulness of the model to galaxy RC data.
The average mass-to-light ratios from RCFM fits yield bulge mass-to-light ratios which are within the error estimates on population synthesis models (PSM). The RCFM disk mass-to-light ratios are approximately a factor of two larger than PSM, consistent with the artifact introduced from the spherical integration technique, see Sec. 2.5 for a full explanation. This artifact can be removed by including an exponential disk integration at the cost of additional computation.
@@ 4.3 Comparing Milky Way models
RCFM fits, as currently formulated, require a choice of a static Milky Way baryon distribution. In Table 9 and LABEL:TSet we compare the RCFM fit results for the two different Milky Way model assumptions used in this paper. As can be seen in the table, for the 36 most reliable galaxies in the SPARC sample the two Milky Way models are equivalent (see Sec. 4.6). However, for the whole SPARC sample of 175 galaxies the RCFM fits using the Sofue-Xue-Jiao Milky Way performs better with an average , versus those from McGaugh-Jiao Milky Way . Based on this, values in Table LABEL:table:M2Light and figures are from RCFM fits assuming the Sofue-Xue-Jiao Milky Way.
The residuals of fits to the SPARC sample are also used to compare the two different MW models (see Sec. 3.2). Histograms of residuals normalized by the error in velocity observations are shown in Fig. 4. In all cases, residuals of model fits to observed velocity data followed a narrow distribution centered at zero with a range of standard deviations in error, albeit with heavy tail features. The behavior of the residuals did not vary greatly between MW models, suggesting that the fitting parameters in our model are robust with respect to differing MW model assumptions at this present level of analysis. With a finer grained analysis and a larger sample, the degeneracy between MW models may be resolved.
A Gaussian fit on residuals from fits using the McGaugh-Jiao MW model gave a mean of , standard deviation of in units of standard deviations in velocity error; whereas a Gaussian fit to the residuals from fits using the Sofue-Xue-Jiao MW model gave a mean of , standard deviation of in those same units. The residuals are shown in Fig. 4. The small values associated with these quantities in both cases provide confidence that our fits match data closely.
@@ 4.4.1 NGC 2841
NGC 2841 is a star dominated, flocculent spiral galaxy, which historically has been regarded as a problematic case for MOND ((Gentile et al., 2011)). RAR finds a good fit for this galaxy after adjusting the Cepheid distance of Mpc by , to Mpc. Cepheid variable stars are the most accurate distance indicators currently available. The RCFM fit is shown in Fig. 6 at the Cepheid based distance, and fit results are compared to dark matter and RAR in Table 2.
| Model | distance(Mpc) | ||||
| Dark Matter Iso | 0.60 | 0.66 | 1.58 | 14.1 | |
| RAR | 0.81 | 0.93 | 1.52 | 15.5 | |
@@ 4.4.2 IC 2574
IC 2574 is a dwarf spiral galaxy, dominated by gas with no central stellar bulge. RAR finds a good fit to this galaxy but adjusts the distance and inclination by to . This galaxy is problematic for dark matter model fits, as the model overestimates the inner RC out to 10 kpc (Navarro et al., 2017). RCFM fits this galaxy successfully, at the reported tip of the red giant branch distance of Mpc, Fig.7, model comparisons are in Table 3.
| Model | ||||
| Dark Matter Iso | 0.77 | – | 2.51 | |
| RAR | 0.07 | – | 1.44 | |
@@ 4.4.3 NGC 3198
NGC 3198 is a barred spiral in Ursa Major, which has been well studied. It has been considered a problem galaxy for MOND, when the distance is a free parameter in fits (Gentile et al., 2011), as the preferred MOND distance is 2 different from that reported from Cepheids. However, in the SPARC database the RAR fit to this galaxy at the Cepheid distance reproduces the RC well. This galaxy has no bulge component, as indicated in Table 4. See Fig. 8 for the RCFM fit.
| Model | distance (MPC) | ||||
| Dark Matter Iso | 0.52 | – | 1.31 | 13.7 | |
| RAR | 0.77 | – | 2.06 | 13.8 | |
@@ 4.4.4 NGC 7814 and NGC 891
| Model | ||||
| Dark Matter Iso | 0.68 | 0.71 | 0.39 | |
| RAR | 1.17 | 0.52 | 1.334 | |
@@
| Model | ||||
| Dark Matter Iso | 0.77 | 1.63 | 1.30 | |
| RAR | 0.32 | — | 25.16 | |
@@ 4.4.5 NGC 5055
| Model | ||||
| Dark matter - Isothermal Halo | 0.26 | - | 6.19 | |
| RAR | 0.56 | - | 7.42 | |
@@
The second RC for NGC 5055 is from the THINGS (de Blok et al., 2008a) at a distance of Mpc from the Hubble Flow, including a stellar bulge component. RCFM fits to both sets of RC data can be seen in Fig.10, at reported distances, using the reported luminous mass models. This galaxy serves as an example of the under-constrained nature of galaxy mass modeling (Conroy et al., 2009). Details of the fits are given in Table 10.
| Model | ||||
| Dark Matter Iso | 0.79 | 0.11 | 8.13 | |
| Dark Matter NFW | 0.79 | 0.11 | 10.31 | |
| RAR | 0.43 | 0.46 | 2.63 | |
@@
4.5 Galaxies for which model fits fail
Quality flags are assigned in the SPARC galaxy database using the following schema; are galaxies with high-quality rotation curves; are galaxies with minor asymmetries and/or rotation curves of lower quality; Q = 3 for galaxies with major asymmetries, strong non-circular motions, and/or offsets between HI and stellar distributions, so that galaxies are not suited to detailed dynamical studies. Of the three models compared in this paper against the 175 SPARC galaxies; RAR fits all galaxies, the RCFM fits 172 of 175 galaxies (fits fail for UGC06628 (a galaxy), and F561-1, UGC04305 ( galaxies)), and the isothermal dark matter model fits 165 of 175 galaxies (fits fail for galaxies D512-2, UGC00634, NGC6789, UGC00891, UGC02023, UGC05999, UGC07232 and UGC09992, and galaxies F567-2 and F574-2).
@@ 4.6 Free Parameter Correlation
To test a physical interpretation of the model’s free parameter with respect to a ratio of total luminosity by the half-light radius (effective radius), we select a subset of galaxies from the SPARC dataset by the following criteria:
@@
Select galaxies with the most accurate and reliable distance estimates (tip of the red giant branch and Cepheid variable stars), rejecting all other galaxies.
@@
Exclude galaxies with quality factor as not suited to dynamical studies due to asymmetries, non-circular motions, and/or offsets between stars and gas, using assigned SPARC factors as in (Lelli et al., 2016).
By this process, a training set of galaxies are selected, as reported in Table 9. We then plot the subset’s values versus the ratio of , and fit the distribution with a power law as shown in Fig. 11. The total luminosity and effective radii numbers used are as reported in the SPARC database. Luminosity measurements are taken in the wavelength of , assuming a solar absolute magnitude of 3.24 at (Oh et al., 2008) in units of solar . As can be seen in in Fig. 11 the subset appears to be highly correlated to the ratio of photometric quantities, though slightly different functionally for the choice of Milky Way.
| Model | ||||
| RAR | 0.46 | 0.74 | 3.89 | |
| RCFM (MW: Sofue-Xue-Jiao) | 1.04 | 0.87 | 1.60 | |
| RCFM (MW: McGaugh-Jiao) | 1.00 | 0.87 | 1.65 | |
@@
A subset of galaxies selected based on best distance estimates and highest quality rotation curve data. Reduced values do not reflect galaxies which fail to converge on a fit; for which, the RCFM and RAR have none, and the isothermal dark matter model has two (NGC6789, UGC07232). The average stellar mass-to-light ratios are presented in units of solar mass per solar luminosity . The RCFM spherical simplification used in RCFM fits for computational and formulaic simplicity, can be seen to yield a disk mass-to-light ratio which is approximately a factor of larger than those from RAR and Dark Matter fits. This is a well known artifact of the spherical assumption, and could be removed with a full disk geometry at the cost of additional computations and free-parameters, see Sec. 2.5 for more information.
| MW: | MW: | ||||
| Galaxy | McGaugh-Jiao | Sofue-Xue-Jiao | Distance | Distance | Inc |
| Name | (Mpc) | Method 111 Distance method: 2 = tip of the red giant branch, 3 = Cepheids. Distance and inclination information from (Lelli et al., 2016). | (deg) | ||
@@
5 Conclusions
At this time, there are many discrepancies between cosmology theory and observations which are resolved by dark matter models (Sanders, 2010; Tully et al., 2014; Naidu et al., 2022). In this paper we address only one such discrepancy, the flat-rotation curve problem of spiral galaxies. This choice is due to the clear symmetry presented in the rotation curve data. Other dark matter problems require cosmological model assumptions, beyond the scope of the current paper.
The rotation curve fitting model (RCFM) presented here reproduces the fitting successes of MOND, RAR (Radial Acceleration Relation) and dark matter models on a sample of 175 well studied galaxy rotation curves (Lelli et al., 2016), but does not modify classical physics. Independent of the interpretation of the various quantities appearing in the formulae presented in this paper, this is a one parameter fit to the data. What is more, the RCFM free parameter is highly correlated with a ratio of observable photometric parameters, which makes this the first direct constraint to luminous mass modeling of galaxies from Doppler shifted spectra.
We emphasize that the only input to the RCFM is the luminous mass. In contrast to previous investigations of gravitational redshifts in this context, which used Galilean subtraction (Misner et al., 1973), this approach sets the frame comparison from the center of the galaxies, using Lorentz-type boosts between galaxy frames in radius. In this way, using gravitational redshifts to define frames, we can explain the discrepancy between the “flat-rotation” curves and the Keplerian luminous mass predictions with only one free parameter. Our results are presented in this paper in comparison to dark matter and RAR fit results for the same galaxy data, from the SPARC database of 175 rotation curves (Lelli et al., 2016). RAR has a simpler functional form than MOND, though the same physics paradigm. The recent Gaia DR3 data release Jiao, Yongjun et al. (2023) has shown that the rotation curve of the Milky Way demonstrates a Keplerian decline from - kpc, consistent with the RCFM model paradigm presented here. The upcoming Large Survey of Space and Time at the Vera C. Rubin Observatory (Željko Ivezić et al., 2019) can falsify this model if the Milky Way rotation curve beyond kpc is found to require a dark matter halo.